OpenStudy (anonymous):

given: f(x)= x^3 + 1 and g(x)= sqare root of x. Find. (A.) f(g(x)) (B.) g(f(x)) (C.) f(x+1) (D.) f(x^2)

OpenStudy (anonymous):

f(x)=x^3+1 f(x)=3x^2

OpenStudy (saifoo.khan):

(A.) f(g(x)) g into f. x^3 + 1 \[\Huge \sqrt{x^3}+1\]

OpenStudy (anonymous):

@ushef i dont get that..

OpenStudy (anonymous):

silly me this is rlly easy! haha

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

are the answers to a and b the same?

OpenStudy (akshay_budhkar):

no answers are differnent..

OpenStudy (anonymous):

You just have to replace x by what's into brackets : \[e.g. f(g(x))=g(x)^{3}+1=\sqrt{x}^{3}+1=x \sqrt{x}+1\]

OpenStudy (anonymous):

\[g(f(x))=\sqrt{f(x)}=\sqrt{x+1}\]...